### v1

# Evaluating LLM-driven User-Intent Formalization for Verification-Aware Languages

Abstract
Verification-aware programming languages such as Dafny and F* provide means to formally specify and prove properties of a program. Although the problem of checking an implementation against a specification can be defined mechanically, there is no algorithmic way of ensuring the correctness of the user-intent formalization for programs — that a specification adheres to the user’s intent behind the program. This is because intent or requirement is expressed informally in natural language and the specification is a formal artefact. However, the advent of large language models (LLMs) has made tremendous strides bridging the gap between informal intent and formal program implementations in the last couple of years, driven in large parts due to benchmarks and automated metrics to evaluate different techniques.
Recent work has developed a framework for evaluating and benchmarking the user-intent formalization problem for mainstream programming languages [12]. However, as we argue in this paper, such an approach does not readily extend to verification-aware languages that support rich specifications (using quantifiers and ghost variables) that cannot be evaluated through dynamic execution. Previous work also required generating program mutants using LLMs to create the benchmark. We advocate an alternate, perhaps simpler approach of symbolically testing specifications to provide an intuitive metric for evaluating the quality of specifications that can be easily instantiated with most verification-aware languages. We demonstrate that our automated metric agrees closely with a GPT-4 generated and human-labeled dataset of roughly 150 Dafny specifications for the popular MBPP code-generation benchmark, yet demonstrates cases where the human labeling is not perfect. We believe our work provides a stepping stone to enable the establishment of a benchmark and research agenda for the problem of user-intent formalization for programs.
Index Terms:
[…]
I Introduction
Formal verification is only as good as the specification it verifies. A formal specification unambiguously defines the (possibly partial) intent behind a program, often in an declarative manner. Although there has been decades of research in advancing the state-of-the-art in automating the problem of verifying an implementation against a specification, relatively less attention has been spent on how to aid the generation and evaluation of specifications or formal requirements. It is well-known that the lack of formal specifications is a significant impediment to deployment of formal verification in production code [10]. On the other hand, although a software is often accompanied by informal intent expressed in natural language comments and API documentation, these intents are seldom formally enforced on the underlying implementation.
Large language models (LLMs) have recently demonstrated potential to bridge the gap between informal intent and formal artefact such as code [9, 5], by performing code generation from natural language. Such progress has been spurred in large parts through the creation of crowd-sourced benchmarks such as HumanEval [9] and Mostly Basic Python Programs (MBPP) [5] with automated program-semantics based metrics (such as tests), unlike NLP metrics such as BLEU scores [24]. These benchmarks allow a community to evaluate their models and techniques on a common set of benchmarks automatically and measure progress. One can hope that the establishment of similar benchmarks and metrics for specification generation (decoupled from code generation) can enable synthesizing useful specifications from informal intent in practice.
Motivated by such a need, Endres et al. [12] describe the problem of generating formal declarative specifications (namely, method postconditions) from informal intent using LLMs and automated metrics for evaluating them — we refer to this as user-intent formalization problem in this paper. For code generation, the quality of an implementation generated from informal intent is measured through the set of hidden validation tests (that the code generator does not have access to). For user-intent formalization, Endres et al. re-purpose code-generation benchmarks (such as HumanEval) and introduce two metrics (a) correctness and (b) completeness, with respect to the set of (hidden) validation tests and the (hidden) reference implementation. Correctness captures that the specification satisfies the reference implementation for all the tests, and completeness captures the power of the specification to discriminate against buggy mutations of the reference implementation (given the set of tests). Further, these mutants were generated through LLMs and are grouped by the subset of tests that fail them. They demonstrate that these automatic metrics closely resemble the quality of specifications as determined through manual analysis. In this paper, we investigate creating benchmarks (and associated metrics) for user-intent formalization for verification-aware languages such as F* [26], Dafny [21] and Verus [20] etc. A verification-aware language supports a rich program logic for expressing specifications and offers the ability to verify them statically using automated theorem provers. However, for most non-trivial programs, the verification requires manually decomposing the problem through the use of ghost variables, intermediate lemmas, invariants and assertions. We focus on the problem of synthesizing and evaluating specifications automatically.
I-A Running example
Consider a snippet of the JSON specification of an example from the MBPP dataset for code generation from natural language in Python [9].
The dataset comes with a set of problems, each containing a natural language prompt prompt, the reference code code as well a set of 3 tests in test_list. Misu et al. [22] port these requirements to Dafny (call it MBPP-DFY), including slight change to the prompt, explicitly representing the Dafny method signature (with a slightly changed name), as well as the test cases [1]. 111We take the snapshot of the repository at COMMIT a57ce24.
Misu et al. use GPT-4 [4] and other language models (with sophisticated prompting) to generate a pair of Dafny specification and implementation, and retain generations where the generated specification is provable for the generated implementation. Below we show the specification component of the GPT-4 generated Dafny artefact for this example (with slightly altered method name and signature over JSON) [3].The two postconditions (marked with ensures) define the specification and use an auxiliary predicate InArray.
Although the implementation satisfies the specification, there is no evidence that the specification indeed captures what the user intended (and specified implicitly in the hidden test cases). The authors perform manual review of the specifications and mark them as either to denote if the specification is respectively, inconsistent, weakly or strongly consistent with the intent expressed in the tests [2]. Therefore, although this work addresses generating verified Dafny programs from informal intent, it relies completely on a human to ensure that the specification matches the intent. In particular, the framework allows for the Dafny code to only satisfy a vacuous specification (such as ensures true) or worse, an incorrect one. This aspect renders this dataset not suited for a benchmark for code generation in verification-aware languages, as any alternate specification will need to be labeled by a human, which makes it subjective as well as non-automated.
I-B Proposal: Symbolically testing specifications
Our objective in this work is to define metrics that can be automatically evaluated to determine the quality of a specification. At the same time, we would also like to establish that these metrics correspond well to what a manual labeling would establish. Further, we decouple the problem of specification generation from code generation, to be able to (a) use the specifications to find bugs in underlying or generated code [12], or (b) use the validated specifications to refine the prompts for code generation.
We first argue that the none of the prior approaches are readily applicable to yield an automatic metric: First, running the tests on the implementation without specifications will not rule out vacuous specifications. Second, one cannot apply the approach of Endres et al. for rich specifications that contain ghost state or constructs that cannot be evaluated at runtime (such as a universal/existential quantifier). One can ignore tests and attempt to statically verify the synthesized specification directly against the hidden reference implementation. This can still allow vacuous specifications such as true. Further, it is unlikely that such verification would be automatic for non-trivial specifications given the need to infer intermediate lemmas and invariants. One can instead provide a hidden reference specification and check the candidate specification for semantic equivalence. However, such a metric would be too strict as it would not be able to distinguish weak and vacuous specifications (such as true) from strong (yet incomplete) specifications.
Next, we propose a method for evaluating user-intent formalization for verification-aware languages that leverages the validation tests and symbolic verification capabilities of these languages. Given a set of tests denoting the a set of consistent input-output pairs, we define the correctness and completeness metrics purely over the tests, without the need for an implementation.
Consider a method with signature m(x):y denoting the name m, input parameters x and output parameters y, a candidate postcondition/summary specification (x, y) and a set of input-output tests T.
I-B1 Correctness
A postcondition (x, y) is correct (or sound) with respect to T if it is consistent with all the input-output pairs in T. In other words, for each the following Hoare-triple [15] holds.
I-B2 Completeness
The completeness measure for a specification given T is the fraction of output mutations of the tests in T that is inconsistent with. Let be a finite set of mutations of T. In this work, we restrict the set of mutants per input to a fixed number (5 for this paper). Let be the largest subset such that for each , the following Hoare-triple does not hold:
Then the completeness measure of with respect to T is (this is inspired by kill-set in mutation-testing literature [16]). For curious readers, we also contrast these with a plausible proposal in Appendix B.
Let us demonstrate an implementation of these Hoare triples as Dafny programs for our running example. For correctness against the first test (in JSON), we create the following Dafny program by providing the body of SharedElements:
The Dafny program symbolically checks (using SMT solvers [6]) that the specification holds for the specific input and output. The two assume statements constrain the content of input parameters a and b to equal values in a1 and a2 respectively. The redundant asserts are needed for the verifier to trigger the quantifiers used in InArray to enable the proof. Finally, the output variable is assigned one of the expected values ([5,4] is also an acceptable value).
The above verification only proves that the specification is correct for the test; a vacuous specification ensures true would also be verified. For completeness, we mutate the output value in result in several ways and check that the verification fails. In this case, mutating the value in result to [6] would fail the first postcondition, since 6 is not present in both the input arrays.
For this program, the above specification is marked as strong_spec by the authors of MBPP-DFY. However, our automated test harness coupled with mutation discovers that the specification is incomplete (score 0.6). When we check the above program with either of the two mutations result := [4] or result := [5], the specification still verifies. This is because the implication ==> in the first postcondition only checks that values in result is present in both the arrays; it does not check that all such common values are present in result! Thus our automated metric is able to assign the above specification a lower score than the full functional specification for this example where the ==> is replaced by <==>.
Note
One may caution against using the verification failure (as used for the completeness checks) as a means to show that the corresponding Hoare-triple does not hold. In other words, given that deductive verifiers are sound, but not necessarily complete, a verification failure is typically interpreted as an unknown outcome. However, note that we check for completeness for a specification against an input-output example only if is proved correct for all the tests in T (including the original test ). Since the two verification conditions differ only in a concrete value (between and ), we conjecture that the underlying reasoning (including quantifiers instantiated) would be quite identical in most cases, and the verification failure strongly indicates that the Hoare-triple does not hold.
II Implementation and evaluation
We report ongoing work in implementing the two metrics for the MBPP-DFY dataset mentioned earlier in Section I-A. Our tool (a 400 line Python script) consumes the method signature, test cases and the candidate specification from the JSON and Dafny files, and creates Dafny programs for verifying the correctness and completeness for each test. It then invokes the Dafny verifier, and reports the aggregate correctness as well completeness score for each specification, averaged over the different test cases; it finally compares the metrics against the labels provided by authors. Of the 153 problems with specifications (a subset written by humans), we have currently managed to apply our tool to check 64 of the specifications.
II-A Mutating values
Each of these examples contain 3 test cases in Dafny format, and we consider up to 5 distinct mutants of the output values for each test. We currently have a simple mutation scheme for the output values. For Booleans, we flip the value between true and false. For integers, we choose a random value between 1 and 10, and randomly add or subtract it. For strings, we choose a random character and either replace one of the characters or append it to the string. For arrays (we only restrict to integer arrays and sequences), we choose between dropping an element or inserting a random value at a randomly chosen index. For our running example of SharedElements, this allows us to create a mutant test case with the array value that demonstrates that the GPT-4 generated specification (marked as strong_spec) is not the most precise specification for this problem.
II-B Results
We briefly report some details of the evaluation, and outline the challenges to handle the remaining examples.
We find that for large majority of examples, the manual labels align with the metrics computed by our tool. That is, for the incorrect specifications, our tool reports a verification failure for correctness check, and report a high completeness score (usually ) for precise specifications. This provides evidence that the automated metrics serve as a good proxy for the user label. One such incorrect specification generated by GPT-4 (and correctly labeled) is the problem of RemoveDuplicates (”task_id” 572) (also the motivating example by Endres et al. [12] demonstrating ambiguity of natural language) where the task is to remove all elements with duplicates, but GPT-4 interprets it as the problem of retaining a single copy of each value. Similarly, for countSubstrings (”task_id” 61), that counts the number of substrings whose sum equals their length, our completeness check gives a low score to the specification (correctly labeled weak_spec) that only ensures count is a non-negative number.
In addition to the SharedElements (”task_id” 2), we found at least 2 more cases where the specification (labeled strong_spec) is weaker than the most precise one. These include the cases for maxAbsDiff (”task_id” 145), and removeElements (”task_id” 161). maxAbsDiff is expected to compute the maximum difference between any two elements in the list; the provided specification only ensures that the result upper bounds the difference between any two elements, but is not the precise bound. removeElements is expected to remove the elements in one array from another; the specification only ensures that any element in the result array is present in the first array and not the second but fails to ensure that the result array is the precise difference (can be ensured by changing ==> by <==>).
Interestingly, we also discovered a few cases where the correctness check failed for specifications labeled strong_spec. These include examples with ”task_id”s 234, 240, and 445. For these examples, the authors had accidentally introduced bugs while while copying from the test cases in Python. For instance, for ”task_id” 234 that computes the cube of a number, the authors mistakenly copied the value 25 instead of 125 for the cube of 5. We noticed that a couple of these errors have been fixed in the latest commit at the time of writing (fb4f53e), but a few still remain (e.g., 445 where the value 32 is replaced by 31).
Our preliminary results demonstrates that the automatic metrics not only helps to add objectivity to the manual labeling, argues for having a quantitative metric for completeness (instead of a Boolean weak_spec vs strong_spec), but also clearly demonstrates the potential to serve as a metric for evaluating specifications for a benchmark.
II-C Limitations
Our implementation currently has a few limitations that precludes analyzing all the 153 specifications correctly. A large class of them stem from inaccurate parsing through a simple regex based parsing of the Dafny files. We expect these (e.g., handling of 2D arrays) to be addressed as we improve the parsing through Dafny AST-based analysis. However, a few fundamental challenges (related to failure to perform proofs of correct specifications) remain that requires further investigation. We note a couple of them here:
Recursive predicates.
Specifications for tasks such as 105 (counting number of true Booleans in an array) require the use of recursive functions countTo that expresses the count of an array in terms of countTo of the tail of the array. Dafny uses a notion of ”fuel” to control the level of unfolding of such predicates, which suffices for inductive proofs. However, for concrete arrays in test cases, these recursive functions need to be unrolled proportional to the length of the array.
Quantifier instantiation.
Specifications for tasks such such as 3 (that classifies a number as a non-prime) uses existential quantifiers. For example, checking a number as non-prime requires checking for the existence of a smaller number that divides . Unlike their usage in inductive proofs where a loop establishes the inductive hypothesis to only instantiate the quantifier on a few expressions, testing a symbolic specification against a concrete value is sometimes non-trivial. For example, to show that the value 97 is a prime, the quantifier needs to be instantiated on all values up to 96.
We have already automated a few failed proofs due to quantifier instantiation by adding intermediate redundant assertions equating two arrays on every index (see assert statements in the body of SharedElements in Sec I-A). We are currently automating specifying the amount of fuel and quantifier instantiation witnesses when ranging over the indices of an array.
III Related work and conclusion
Although the literature on specification mining is fairly rich, it has historically focused on techniques that infer specifications of an implementation through dynamic [13] and recently neural techniques [25]. The objective of these techniques is to learn invariants from a few runs that generalize to unseen executions. Static [14, 19, 23] and more recently neural approaches [27, 17] have also been applied to the problem (inductive) invariant generation to aid program proofs. On the other hand, work on user-intent formalization (i.e., translating natural language comments to specifications) [7, 12] synthesize the intended specification, even in the absence of an implementation. The problem of generating test oracle assertions have also been investigated [11], but they only apply to single input or test prefix. TiCoder [18] introduced the term ”user-intent formalization” over test cases for interactive code generation, and alluded to metrics for measuring quality of weak specifications (namely, tests) in terms of correctness (user acceptance) and completeness (prunes buggy codes). The work on autoformalization [28] for translating natural language comments to mathematical theorems is closest to our work. However, these techniques do not use semantic checks (such as tests and symbolic verification) to ensure the quality of the generated formal mathematical statements; instead they leverage textual similarity metrics such as BLEU.
Based on the encouraging results so far, we are optimistic to create a benchmark with the metrics with a large fraction of examples from MBPP-DFY dataset. Just as the creation of MBPP and HumanEval has turbo-charged progress on high-quality code generation, we hope that the benchmark will accelerate research and practice of synthesizing high-quality specifications from informal intent in verification-aware languages. This when coupled with work on program and proof synthesis (given a specification) can greatly lower the cost required to create formally verified modules in the future [8].
Appendix A Extended version of the JSON examples
Appendix B Alternate proposal
A reader may wonder if the following check achieves a similar objective:
There are two issues with this formulation: (a) this only provides a Boolean metric that only rewards a precise specification that constrains y to a unique value . For example, it would award the strong (yet imprecise) specifications of SharedElements to 0. Secondly, (b) this formulation is also too strong when the precise functional specification allows for non-determinism in the output. Consider the running example of SharedElements, where the specification allows for the output to be one of either for the given inputs.